OpenStudy (anonymous):

Write an equation of a line that passes through the point (7, 3) and is parallel to the line y = negative 2 over 3x + 3.

OpenStudy (anonymous):

A. y = 3 over 2x − 3 B. y = 3 over 2x + 3 C. y = negative 2 over 3x + 23 over 3 D. y = negative 2 over 3x − 23 over 3

OpenStudy (anonymous):

@jim_thompson5910 @mathmate

OpenStudy (laughoutloud):

ok so we are going to use point slope form

OpenStudy (laughoutloud):

do you know the point slope formula

OpenStudy (laughoutloud):

@cheetah21

OpenStudy (anonymous):

is y2 -y1/x2-x1?

OpenStudy (laughoutloud):

no that is the formula for finding slope the formula for point slope is \[y -y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

Oh okay my bad so how do we do that ?

OpenStudy (laughoutloud):

well they pretty much give you everything what is the slope?

OpenStudy (laughoutloud):

*hint since it is parallel the slope are the same

OpenStudy (anonymous):

the slope is 2?

OpenStudy (laughoutloud):

is this how your equation looks \[y=\frac{ 2 }{ 3 }x+3\] is so then the slope is 2/3x

OpenStudy (anonymous):

yeah that how it looks except the fraction is negative

OpenStudy (laughoutloud):

oh okay sorry so the slope is -2/3x

OpenStudy (anonymous):

its okay(: soo whats next

OpenStudy (laughoutloud):

now you plug in the values the slope for m and the point (7,3) for x(1) and y(1)

OpenStudy (anonymous):

so y - 3 = m(x-7) ?

OpenStudy (laughoutloud):

yes but plug in the slope for m

OpenStudy (anonymous):

Okay i changed the m to -2/3, now what cause im having a little trouble solving it

OpenStudy (laughoutloud):

solve for y

OpenStudy (anonymous):

7y times -2/3?

OpenStudy (laughoutloud):

k idk what you did to get that lol but when you distribute you should get \[y-3=-\frac{ 2 }{ 3 }x-\frac{ 14 }{ 3 }\]

OpenStudy (anonymous):

I got that i just dont know where to go from there

OpenStudy (laughoutloud):

wait it should be \[y-3=-\frac{ 2 }{ 3 }+\frac{ 14 }{ 3 }\]

OpenStudy (laughoutloud):

then combine like terms

OpenStudy (laughoutloud):



OpenStudy (anonymous):

y-3=16/3

OpenStudy (laughoutloud):